Bibliography of N. C. Phillips

March 2001

1. N. C. Phillips, K-Theoretic Freeness of Actions of Finite Groups on C*- Algebras, Ph.D. Thesis, University of California at Berkeley, 1984.

2. N. C. Phillips, K-theoretic freeness of finite group actions on C*-algebras, pages 227–243 in: Group Actions on Rings, S. Montgomery (ed.), Contemporary Mathematics vol. 43, Amer. Math. Soc., Providence RI, 1985.

3. N. C. Phillips, Equivariant K-Theory and Freeness of Group Actions on C*-Algebras, Springer-Verlag Lecture Notes in Mathematics No. 1274, Springer-Verlag, Berlin, Heidelberg, New York, London, Paris, Tokyo, 1987.

4. N. C. Phillips, Equivariant K-theory for proper actions and C*-algebras, pages 175–204 in: Index Theory of Elliptic Operators, Foliations, and Operator Algebras, J. Kaminker, K. C. Millett, and C. Schochet (eds.), Contemporary Mathematics vol. 70, Amer. Math. Soc., Providence RI, 1988.

5. N. C. Phillips, Equivariant K-theory for proper actions II: Some cases in which finite dimensional bundles suffice, pages 205–227 in: Index Theory of Elliptic Operators, Foliations, and Operator Algebras, J. Kaminker, K. C. Millett, and C. Schochet (eds.), Contemporary Mathematics vol. 70, Amer. Math. Soc., Providence RI, 1988.

6. N. C. Phillips, Inverse limits of C*-algebras, J. Operator Theory 19(1988), 159–195.

7. N. C. Phillips, A new approach to the multipliers of Pedersen’s ideal, Proc. Amer. Math. Soc. 104(1988), 861–867.

8. N. C. Phillips, Inverse limits of C*-algebras and applications, pages 127– 186 in: Operator Algebras and Applications Volume 1: Structure Theory: K-Theory, Geometry and Topology, London Math. Soc. Lecture Note Series No. 135, Cambridge University Press, Cambridge, New York, New Rochelle, Melbourne, Sydney, 1988.

9. N. C. Phillips, Equivariant K-Theory for Proper Actions, Pitman Research Notes in Mathematics No. 178, Longman Scientific and Technical, Essex, England, 1989.

10. N. C. Phillips, Representable K-theory for σ-C*-algebras, K-Theory 3(1989), 441–478.

11. N. C. Phillips, The Atiyah-Segal completion theorem for C*-algebras, K-Theory 3(1989), 479–504.

12. N. C. Phillips, Classifying algebras for the K-theory of σ-C*-algebras, Canadian J. Math. 41(1989), 1021–1089.

13. N. C. Phillips, K-theory and noncommutative homotopy theory, Proceedings of Symposia in Pure Mathematics vol. 51 (1990), part 2, 255–265.

14. N. C. Phillips, Equivariant K-theory for proper actions III: uniform bundles and discrete groups, preprint (1990).

15. N. C. Phillips, A direct prof of the Bott periodicity theorem in noncommutative homotopy theory, preprint (1990).

16. N. C. Phillips, C^{∞} loop algebras and noncommutative Bott periodicity, Trans. Amer. Math. Soc. 325(1991), 631–651.

17. N. C. Phillips, K-theory for Fréchet algebras, International J. Math. 2(1991), 77–129.

18. N. C. Phillips, Simple C*-algebras with the property weak (FU), Math. Scand. 69(1991), 127–151.

19. N. C. Phillips and J. R. Ringrose, Exponential rank in operator algebras, pages 395–413 in: Current Topics in Operator Algebras, H. Araki, etc. (eds.), World Scientific, Singapore, New Jersey, London, Hong Kong, 1991.

20. N. C. Phillips, Five problems on operator algebras, pages 133–138 in: Selfadjoint and Nonselfadjoint Operator Algebras and Operator Theory, Robert S. Doran (ed.), Contemporary Mathematics vol. 120, Amer. Math. Soc., Providence RI, 1991.

21. N. C. Phillips, The Toeplitz operator proof of noncommutative Bott periodicity, J. Austral. Math. Soc. Ser. A 53(1992), 229–251.

22. N. C. Phillips, The rectifiable metric on the space of projections in a C*- algebra, International J. Math. 3(1992), 679–698.

23. N. C. Phillips, Approximation by unitaries with finite spectrum in purely infinite simple C*-algebras, J. Funct. Anal. 120(1994), 98–106.

24. N. C. Phillips, How many exponentials?, Amer. J. Math. 116(1994), 1513– 1543.

25. N. C. Phillips, The C*-projective length of n-homogeneous C*-algebras, J. Operator Theory 31(1994), 253–276.

26. N. C. Phillips, Reduction of exponential rank in direct limits of C*-algebras, Canadian J. Math. 46(1994), 818–853.

27. N. C. Phillips and L. B. Schweitzer, Representable K-theory of smooth crossed products by R and Z, Trans. Amer. Math. Soc. 344(1994), 173– 201.

28. N. C. Phillips, A survey of exponential rank, pages 353–399 in: C*- Algebras: 1943–1993 (A Fifty Year Celebration), Robert S. Doran (ed.), Contemporary Mathematics vol. 167, Amer. Math. Soc., Providence RI, 1994.

29. N. C. Phillips, Exponential length and traces, Proc. Royal Soc. Edinburgh (Section A), 125(1995), 13–29.

30. N. C. Phillips, Inner derivations on σ-C*-algebras, Mathematische Nachrichten 176(1995), 243–247.

31. N. C. Phillips, Every invertible Hilbert space operator is the product of seven positive operators, Canadian Math. Bull. (ser. 2) 38(1995), 230– 236.

32. H. Lin and N. C. Phillips, Classification of direct limits of even Cuntzcircle algebras, Memoirs Amer. Math. Soc. no. 565 (1995).

33. H. Lin and N. C. Phillips, Approximate unitary equivalence of homomorphisms from O_{∞}, J. reine ang. Math. 464(1995), 173–186.

34. H. Lin and N. C. Phillips, Almost multiplicative morphisms and the Cuntz algebra O_{2}, International J. Math. 6(1995), 625–643.

35. N. C. Phillips, Factorization problems in the invertible group of a homogeneous C*-algebra, Pacific J. Math. 174(1996), 215–246.

36. G. J. Murphy and N. C. Phillips, C*-algebras with the approximate positive factorization property, Trans. Amer. Math. Soc. 348(1996), 2291–2306.

37. N. C. Phillips, UHF algebras have neither property (FS′) nor exponential rank 1 , Bull. Sci. Math. 121(1997), 549–560.

38. N. C. Phillips, Approximate unitary equivalence of homomorphisms from odd Cuntz algebras, Fields Institute Communications 13(1997), 243–255.

39. N. C. Phillips, Continuous embedding of the rotation algebras in the Cuntz algebra O_{2}, pages 65–73 in: Operator Algebras and Quantum Field Theory, S. Doplicher, R. Longo, J. E. Roberts, L. Zsido (eds.), International Press, Cambridge MA, 1997. (Reports results of computer experiments.)

40. N. C. Phillips and N. Weaver, Modules with norms which take values in a C*-algebra, Pacific J. Math. 185(1998), 163–181.

41. Q. Lin and N. C. Phillips, Ordered K-theory for C*-algebras of minimal homeomorphisms, pages 289–314 in: Operator Algebras and Operator Theory, L. Ge, etc. (eds.), Contemporary Mathematics vol. 228, 1998.

42. E. Kirchberg and N. C. Phillips, Embedding of exact C*-algebras in the Cuntz algebra O_{2}, J. reine ang. Math. 525(2000), 17–53.

43. E. Kirchberg and N. C. Phillips, Embedding of continuous fields of C*- algebras in the Cuntz algebra O_{2}, J. reine ang. Math. 525(2000), 55–94.

44. N. C. Phillips, A classification theorem for nuclear purely infinite simple C*-algebras, Doc. Math. 5(2000), 49–114 (electronic).

45. L. Ge and N. C. Phillips, Nonisomorphic simple exact C*-algebras with the same Elliott and Haagerup invariants, preprint.

46. N. C. Phillips, Recursive subhomogeneous algebras, preprint.

47. N. C. Phillips, Cancellation and stable rank for direct limits of recursive subhomogeneous algebras, preprint.

48. N. C. Phillips, Real and exponential rank of tensor products with O_{∞}, J. Operator Theory, to appear.

49. N. C. Phillips, Type 1 C*-algebras not isomorphic to their opposite algebras, International J. Math., to appear.

50. M. Nagisa, H. Osaka, and N. C. Phillips, Ranks of algebras of continuous C*-algebra valued functions, Canadian J. Math., to appear.

51. Q. Lin and N. C. Phillips, Direct limit decomposition for C*-algebras of minimal diffeomorphisms, preprint.

52. H. Lin, N. C. Phillips, and S. Zhang, Connectedness of the unitary group of certain multiplier algebras, in preparation. (The proofs are all done.)

53. Q. Lin and N. C. Phillips, Generic Rokhlin towers in smooth minimal dynamical systems, in preparation (almost done).

54. Q. Lin and N. C. Phillips, The structure of C*-algebras of minimal diffeomorphisms, in preparation.

55. N. C. Phillips, Many nonisomorphic nonnuclear simple C*-algebras with the same Elliott invariant, in preparation. (The proofs are all done.)